# Social Choice for AI Alignment: Dealing with Diverse Human Feedback

### v1

@@ 3 What are the Collective Decision Problems and their Alternatives in this Context?
If we want to use methods from social choice for the purpose of aligning AI systems, we first need to specify what the concrete options/objects are, before we can start collecting preferences over them and make actual or simulated collective choices between them. These options are called alternatives in social choice theory. In some contexts, the set of alternatives is easy to comprehend and enumerate, as when the alternatives are candidates for a position or an award. In other settings, there are exponentially many alternatives, but the set is still easy to comprehend, e.g., when there are $n$ propositions and each of them must be either accepted or rejected (Lang, 2007).
@@
When considering the alignment of AI systems, it is harder to see exactly how best to think about the relevant set of alternatives for evaluation. In principle, it could be the set of all AI systems or all possible parameterizations of a given network architecture, but this would surely be conceptually intractable.
@@
In the context of an LLM, the RLHF approach traditionally asks the evaluator to choose between a small, explicit set of alternative responses to a single prompt, with each response sampled from the LLM’s output distribution. Alternately, we could consider all possible responses as alternatives. While this response set is too large to explicitly enumerate, the evaluators can still indicate their preference by providing the preferred response themselves. Such exemplars are often used for fine-tuning and can be used to learn evaluators’ preferences and generate responses that well-represent them (Fish et al., 2023). While this does not address questions about how to generalize beyond a single prompt, it is a useful way of conceptualizing the alternatives.
@@
One might conceive of the alternatives as probability distributions over responses. This is natural, as LLMs are typically configured to respond stochastically to a prompt. This might be desirable not only for creativity but also to promote fairness and representativeness of responses. For example, in response to a controversial question, fairness might militate against an LLM always giving the same answer, as any one answer will inevitably omit some relevant considerations on one side of a debate. There is a large literature on social choice rules whose outputs are probability distributions. The inputs to such a rule could be the evaluators’ stated explicit preferences between distributions (Fishburn 1973, Ch. 18), but they could also be stated preferences between plain alternatives (Brandt, 2017). Indeed, the type of objects chosen by a social choice rule (e.g., distributions over responses) need not match the type of objects about which individuals state their preferences or evaluations (e.g., responses). This is important, since probability distributions over large sets of responses may be particularly difficult for evaluators to reliably compare.
@@
A multi-winner rule (Faliszewski et al., 2017; Elkind et al., 2017) could be a middle ground between a deterministic single-winner social choice rule and a probabilistic rule. Such a rule picks a small, predetermined number of answers that best reflect what the voters want. These winning answers could then be combined into a single response that lists these winning answers as bullet points to provide the user with a representative overview of possible answers.
@@ 4 Who Provides the Human Feedback?
Let us assume that there is a population of people, the stakeholder population, who will be affected by an AI system and whose preferences would therefore ideally be taken into account in aligning the AI system. [^3] There may also be stakeholders, such as small children and non-human animals, whose feedback we cannot easily elicit. In that case, we may consider feedback from humans who are charged with representing their interests. Unfortunately, it may be infeasible to elicit feedback from all members of the stakeholder population, so we must select some smaller group from which to elicit feedback. For example, one could try to select a suitably representative subset of the population such that the alignment obtained using feedback from the subset sufficiently approximates the alignment that would be obtained using feedback from the full stakeholder population. Here one could draw on ongoing work in social choice theory on how to select citizens’ assemblies that are representative of a full population (e.g., Flanigan et al. 2021; Landemore & Fourniau 2022), as well as work in statistics on efficient stratified sampling (e.g., Meng 2013).
Another approach would be to allow the full stakeholder population to vote on their representatives in some way. This could be done, for example, with a voting procedure that is designed to elect assemblies that are proportionally representative (see, e.g., Ch. 4 of Lackner & Skowron 2023). Additionally, stakeholders might be allowed to delegate their feedback rights to others (who may in turn delegate, etc.), as in liquid democracy (see Paulin 2020).
As of now, earlier work has used evaluator recruitment methods such as Mechanical Turk (Freedman et al., 2020; Bai et al., 2022a); Upwork, Scale AI, or Lionbridge (Stiennon et al., 2020; Ziegler et al., 2019); and purpose-built platforms (Noothigattu et al., 2018). We believe this component of the RLHF pipeline deserves a more in-depth discussion, including one informed by social choice theory.
@@ 5 What is the Format of Human Feedback?
As we have discussed, human feedback for AI systems can come in various forms; which of these are most natural and useful? Here, we can draw on a significant literature on preference elicitation (see, e.g., Sandholm & Boutilier 2006), studying how best to query agents for their preferences in a variety of domains. [^4] Incidentally, recently, it has been proposed to use LLMs as a tool to help do so (Li et al., 2023). This literature is closely tied to that of communication complexity (e.g., Kushilevitz & Nisan 1997), which is concerned with minimizing the number of bits that need to be communicated to achieve something. (Though the number of bits communicated is of course generally not the perfect way to measure how much effort a human participant needs to expend to answer queries.) Preference elicitation and communication complexity have also been studied in voting settings (Conitzer & Sandholm, 2002, 2005; Service & Adams, 2012).
@@ 5.1 Multiple Format Options
In general, we want the type of input or feedback that we ask of humans to be (1) natural to give, (2) informative about their preferences and values, and (3) of a type that can be used to align AI systems. For example, with current methods, having humans comment on an AI output in an open-ended text box may satisfy 1 and 2, but not 3. Having them sort responses alphabetically may satisfy 1 and 3, but not 2. Having them directly rank neural networks based on inspecting their weights may satisfy 3 but not 1 or 2.
It should be noted that different choices for the type of input or feedback can lead to differently aligned systems, especially if we do not understand the behavioral effects of the different types of input. For example, McElfresh et al. (2021) introduce (in the context of feedback on kidney allocation) an indecision option among the available choices and reject several natural hypotheses about how the resulting data relate to those obtained without that option.
One question is whether we should actually let individual humans choose the format in which they give input or feedback. In traditional social choice, this is uncommon, although there may be some flexibility in how preferences are expressed or input is given (e.g., allowing voters to not give a complete ranking but rather only rank a few alternatives (Halpern et al., 2023), or to give numerical ratings instead of ordinal rankings), as well as some variety in the interaction mechanism to get to that expression of preferences (e.g., one can vote for candidates individually but also pull a lever that corresponds to voting for exactly the candidates of a single party).
It is easy to imagine giving evaluators the choice between a range of different ways to give their input or feedback on various aspects of the system’s behavior or behavioral patterns or rules (e.g., individual responses, whole dialogue sessions, longterm interaction with the same user, or published guiding principles) and various dimensions of desirability, which is emerging as fine-grained RLHF (Wu et al., 2023) or optimizing attributes in the data (Dong et al., 2023), relating to various values such as “truthfulness”, “harmlessness”, “fairness”, etc., and to allow them to give that feedback in various ways: approving/disapproving, making pairwise comparison statements of the form “I like A better than B”, giving full or partial rankings of the form “A is best, B 2nd-best, …”, giving precise or imprecise ratings of the form “I rate A between 7 and 9”, or even by giving free-form verbal feedback that the LLM then interprets and converts into some formal data such as a partial ordering. This heterogeneous data could then be transformed in some formal way into a common, sufficiently expressive data structure, such as a utility function.
5.2 Dealing with Diverse and Informal Feedback
@@
Recall that in RLHF, human feedback is typically used to train a reward (or “preference”) model whose job it is to map any possible AI system response to a numerical rating. The concept of reward models could also be used to convert the diverse input or feedback of a single evaluator into a common form, in order to then aggregate it with the input of other evaluators to steer an AI system.
@@
First, an individual evaluation interpretation model $\phi$ could be trained to map a tuple of inputs of the form $(x,\mathcal{Y},f_{i},e,y)$ to a numerical evaluation $r$. As before, $x$ represents a prompt to the AI system, $\mathcal{Y}$ the set of possible AI responses, and $y\in\mathcal{Y}$ a particular response. Moreover, vector $f_{i}$ represents the relevant features of a certain evaluator $i$, and $e$ shall be a language representation of $i$’s feedback on possible responses $\mathcal{Y}$ to $x$, containing preference- and evaluation-related statements of whatever type (see Section 5.1). In practice, $\phi$ would likely be based on an LLM pretrained to understand the texts $x$, $\mathcal{Y}$, $e$, and $y$, that is then fine-tuned to the interpretation task described above. Then the output $r=\phi(x,\mathcal{Y},f_{i},e,y)$ of $\phi$ is a numerical rating of $y$ given by evaluator $i$ that is trained to be (approximately) consistent with the verbal evaluation $e$ of that evaluator. We note that this task can be seen as a form of meta-learning.
@@
One could then use the trained evaluation interpretation model $\phi$ to train another model—an individual preference model $\psi$—that skips verbal evaluations and directly maps inputs $(x,\mathcal{Y},f_{i},y)$ to ratings $r=\psi(x,\mathcal{Y},f_{i},y)$. Namely, any tuple $(x,\mathcal{Y},f_{i},e)$ can be converted into supervised training data $\big{(}(x,\mathcal{Y},f_{i},y),\phi(x,\mathcal{Y},f_{i},e,y)\big{)}_{y\in\mathcal{Y}}$ for $\psi$, containing simulated ratings $r=\phi(x,\mathcal{Y},f_{i},e,y)$. The hope is that the individual preference model $\psi$ would be able to simulate the rating of any evaluator (represented by their features $f_{i}$), as long as the evaluator, prompt, and response set come from the same distribution as the one $\psi$ was trained on. Similar to the preference models used in current RLHF, $\psi$ could finally be used to fine-tune the actual AI system or steer its behavior in real time. In fact, if the evaluators’ features $f_{i}$ were omitted in the training process sketched above, $\psi$ would be a preference model of the same type as is already used in RLHF and could readily be used for it. This would, however, conflate the evaluations of the (possibly not proportionally representative) set of evaluators used in training in a rather uncontrolled and potentially confusing way. An arguably better way of making use of $\psi$ is, therefore, to indeed make use of evaluators’ features $f_{i}$ in training and add an additional social choice step to the RLHF pipeline or the AI system’s real-time decision-making procedure. Below we sketch several ways in which this might be done.
6 How can Diverse Individual Input or Feedback be Incorporated?
@@
Here we sketch several variants of two approaches for including diverse input or feedback into AI systems in a consistent way using methods from social choice theory. The first suggests adding an additional preference aggregation step somewhere during training, thereby turning RLHF into RLCHF: Reinforcement Learning from Collective Human Feedback. The second approach instead suggests adding an additional simulated collective decision step somewhere in the training or the system’s real-time decision procedure, similar to Bakker et al. (2022) and Jarrett et al. (2023).
@@ 6.1 Proposal: Reinforcement Learning from Collective Human Feedback (RLCHF)
Preference aggregation could be incorporated as an additional step into RLHF in several ways, from early to rather late in the RLHF pipeline. For clarity of exposition, assume a simple version of rankings-based RLHF that (1) takes a database of prompts $x$ together with corresponding sets of possible responses $\mathcal{Y}$, (2) asks one associated evaluator $i(x,\mathcal{Y})$ to provide a ranking $R(x,\mathcal{Y})$ of the elements of $\mathcal{Y}$, (3) turns this ranking into $|\mathcal{Y}|$ many data points for training a common preference model $\varrho$ that produces numerical ratings $r=\varrho(x,y)$, and (4) uses these ratings as rewards in fine-tuning the actual LLM via reinforcement learning.
The earliest point to introduce preference aggregation in this pipeline would be between steps (2) and (3). Instead of a single evaluator $i(x,\mathcal{Y})$, we may ask the members of a jury $J(x,\mathcal{Y})$ of evaluators to provide individual rankings $R_{j}$. Using some ordinal social welfare function $F$, those rankings can then be aggregated into a collective ranking $R=F((R_{j})_{j\in J})$ to use it in step (3). This approach could be termed “RLCHF using aggregated rankings” and is shown in Fig. 2.
@@
Alternatively, one could use cardinal rather than ordinal preference aggregation at a later point in the pipeline: between steps (3) and (4). For this, change step (3) so that not a model of common but of individual preferences is trained, mapping pair $(x,\mathcal{Y})$ and evaluator $i$ with features $f_{i}$ to predicted ratings $r_{i}=\psi(x,f_{i},y)$. Also generate a large collection of feature vectors $f_{1},\dots,f_{N}$ that is representative of the stakeholder population. Then a cardinal social welfare function $W$ can be used to aggregate into one rating $\varrho(x,y)=W(\psi(x,f_{1},y),\dots,\psi(x,f_{N},y))$ which can be used in step (4). This approach could be termed “RLCHF using evaluator features and aggregated ratings” and is shown in Fig. 3.
@@ 6.2 Proposal: Simulated Collective Decisions
RLCHF, as described above, keeps the reinforcement learning step that requires numerical rewards, and it uses ordinal or cardinal preference aggregation to produce these said rewards for all possible responses $y\in\mathcal{Y}$. A different approach would replace reinforcement learning by something else and introduce social choice methods in the form of simulated collective decisions rather than preference aggregation.
For one thing, one could modify “RLCHF using evaluator features and aggregated ratings” into “Supervised Learning from Simulated Collective Decisions”, as shown in Fig. 4. For this, in step (3) from above, use the individual preference model $r_{i}=\psi(x,f_{i},y)$ and feature vectors $f_{1},\dots,f_{N}$ not to produce an aggregated rating but to simulate a collective choice that picks a single winning response $y^{*}=C\big{(}(\psi(x,f_{j},y))_{y\in\mathcal{Y},j=1,\dots,N}\big{)}$. Here, $C$ is now a single-winner social choice function. Then in step (4), use data point $(x,y^{*})$ to train the actual AI system via supervised (rather than reinforcement) learning. Instead of picking a single winner $y^{*}$, we could also use a multi-winner social choice function $C$ that outputs, say, a set of three responses $(y^{\prime},y^{\prime\prime},y^{\prime\prime\prime})$. These can then be (creatively) combined into a single response, for example, by merging them into a bullet-point list and adding a sentence “The following are (three) typical answers to your question: …” at the beginning.
A more radical modification would drop the fine-tuning-via-learning step altogether (leaving the LLM only pretrained) and rather simulate the collective choice at inference time. Whenever the live system is prompted with some $x$, generate $k\gg 1$ many candidate responses $y_{i}$ and $N\gg 1$ many evaluator feature vectors $f_{j}$ representative of the stakeholder population for the problem $(x,\mathcal{Y})$, and directly return the winner $y^{*}=C\big{(}(\psi(x,f_{j},y_{i}))_{j,i=1}^{N,k}\big{)}$ of the simulated collective choice. Here, too, $C$ could be a multi-winner or probabilistic social choice rule.

### v2

@@ 3 What Are the Collective Decision Problems and their Alternatives in this Context?
If we want to use social choice methods for aligning AI systems, we first need to specify what the concrete options/objects are, then collect preferences over them, and finally make actual or simulated collective choices between them. These options are called alternatives. In some contexts, the set of alternatives is easy to comprehend and enumerate, e.g. candidates for a position. In other settings, there are exponentially many alternatives, but the set is still easy to comprehend, e.g., if each of $n$ propositions must be either accepted or rejected (Lang, 2007).
@@
For AI alignment, it is harder to see how best to determine the set of alternatives for evaluation. It could be the set of all possible parameterizations of a given network architecture, but this would surely be conceptually intractable.
@@
For an LLM, most RLHF approaches ask the evaluator to choose between a small, explicit set of alternative responses to some prompt, handwritten or sampled from the pretrained LLM. Alternately, we could consider all possible responses as alternatives. Evaluators could then indicate their preference by providing the preferred response themselves. Such exemplars are often used for fine-tuning and can be used to learn evaluators’ preferences and generate responses that well-represent them (Fish et al., 2023). While this does not address questions about how to generalize beyond a single prompt, it is a useful way of conceptualizing the alternatives.
@@
One might conceive of the alternatives as probability distributions over responses; LLMs are anyway configured to respond stochastically. This might be desirable not only for creativity but also to promote fairness and representativeness of responses. In response to a controversial question, fairness might militate against always giving the same answer, as any one answer will inevitably omit some relevant considerations on one side of a debate. There is a large literature on social choice rules that output probability distributions. Their input could be the evaluators’ stated preferences between distributions (Fishburn 1973, Ch. 18), or stated preferences between plain alternatives (Brandt, 2017), since evaluators might have difficulties comparing probability distributions. Indeed, the type of objects chosen by a social choice rule (e.g., distributions over responses) need not match the type of objects about which individuals state preferences (e.g., responses).
@@
Multi-winner rules (Faliszewski et al., 2017; Elkind et al., 2017) form a middle ground between deterministic single-winner rules and probabilistic rules. They pick a small, predetermined number of answers that best reflect what the voters want, which could be combined into a single response that lists these answers as bullet points to provide the user with a representative overview of possible answers.
@@ 4 Who Provides the Human Feedback?
Let us assume a stakeholder population of people who will be affected by an AI system and whose preferences would therefore ideally be considered in aligning the AI. [^2] Some stakeholders, such as small children and non-human animals, whose preferences we cannot easily elicit might need to be represented by dedicated others. When it is infeasible to elicit feedback from all stakeholders, we must select a smaller group to query. One can try to select a suitably representative subset such that the alignment obtained using their feedback sufficiently approximates the alignment that would be obtained from all stakeholders’ feedback. Here one could draw on ongoing work in social choice theory on how to select citizens’ assemblies that are representative of a full population (e.g., Flanigan et al. 2021; Landemore & Fourniau 2022), as well as work in statistics on efficient stratified sampling (e.g., Meng 2013).
Alternatively, one could let stakeholders vote on their representatives, e.g., with a voting procedure designed to elect proportionally representative assemblies (see, e.g., Ch. 4 of Lackner & Skowron 2023). Stakeholders might also delegate their feedback rights to others (who may in turn delegate, etc.), as in liquid democracy (see Paulin 2020).
Work up to now has used evaluator recruitment methods such as Mechanical Turk (Freedman et al., 2020; Bai et al., 2022a); Upwork, Scale AI, or Lionbridge (Stiennon et al., 2020; Ziegler et al., 2019); and purpose-built platforms (Noothigattu et al., 2018). We believe this component of the RLHF pipeline deserves a more in-depth discussion, including one informed by social choice theory.
@@ 5 What Is the Format of Human Feedback?
As discussed, human feedback for AI systems can have various forms; which of these are most natural and useful? Here, we can draw on a significant literature on preference elicitation (see, e.g., Sandholm & Boutilier 2006), studying how best to query agents for their preferences in several domains. [^3] Incidentally, Li et al. (2023) propose to use LLMs for this. This literature is closely tied to that of communication complexity (e.g., Kushilevitz & Nisan 1997), which is concerned with minimizing the number of bits needed to communicate something. Both of these topics have also been studied in voting settings (Conitzer & Sandholm, 2002, 2005; Service & Adams, 2012).
@@ 5.1 Multiple Format Options
In general, we want the type of input or feedback that we ask of humans to be (1) natural to give, (2) informative about their preferences and values, and (3) of a type that can be used to align AI systems. For example, having humans comment on an AI output in an open-ended text box may satisfy 1 and 2, but not 3 (at least, with current methods). Having them sort responses alphabetically may satisfy 1 and 3, but not 2. Having them directly rank neural networks based on inspecting their weights may satisfy 3 but not 1 or 2. Needless to say, different choices for the type of input may lead to differently aligned systems and have different behavioral effects on humans (cf. Section 8).
Perhaps one should let individuals choose the format in which they give input or feedback. In traditional social choice this is uncommon—though there may be some flexibility in how preferences are expressed (e.g., allowing voters to only rank a few alternatives rather than all (Halpern et al., 2023), or to give numerical ratings instead of rankings) and some variety in the interaction mechanism (e.g., one can vote for candidates individually, or pull a lever to vote for all candidates of a single party at once). Going forward, however, it is easy to imagine giving evaluators choices from a range of different ways to give their input. For example, the input could be individual responses, whole dialogue sessions, longterm interactions with the same user, or published guiding principles. The feedback could be expressed as approval/disapproval votes, pairwise comparisons of alternatives, full or partial rankings of the alternatives, giving precise or imprecise ratings of the form “I rate A between 7 and 9”, or even in free-form verbal feedback that an LLM can interpret as some formal preference data such as a partial ordering. Moreover, the evaluation could be on various aspects of the system’s behavioral patterns, as done in fine-grained RLHF (Wu et al., 2023), or in RLHF that optimizes for multiple attributes (Dong et al., 2023) such as helpfulness, humor, toxicity, etc. Section 5.2 discusses how we may process all this heterogeneous data.
5.2 Processing Diverse and Informal Feedback
@@
Recall that in RLHF, human feedback is typically used to train a reward (or “preference”) model, which maps any possible AI system response to a numerical rating. The concept of reward models could also be used to convert an evaluator’s diverse input into a common form, in order to then aggregate it with other evaluators’ input.
@@
First, an individual evaluation interpretation model $\phi$ could be trained to map a tuple of inputs of the form $(x,\mathcal{Y},f_{i},e,y)$ to a numerical evaluation $r$. As before, $x$ represents a prompt to the AI system, $\mathcal{Y}$ the set of possible AI responses, and $y\in\mathcal{Y}$ a particular response. Vector $f_{i}$ represents the relevant features of a certain evaluator $i$, and $e$ is a language representation of $i$’s feedback on possible responses $\mathcal{Y}$ to $x$, containing preference- and evaluation-related statements of various types (see Section 5.1). In practice, $\phi$ would likely be based on an LLM pretrained to understand the texts $x$, $\mathcal{Y}$, $e$, and $y$, that is then fine-tuned to the interpretation task described above. Then the output $r=\phi(x,\mathcal{Y},f_{i},e,y)$ of $\phi$ is a numerical rating of $y$ given by evaluator $i$ that is trained to be (approximately) consistent with the verbal evaluation $e$ of that evaluator. We note that this task can be seen as a form of meta-learning.
@@
One could then use the trained evaluation interpretation model $\phi$ to train another model—an individual preference model $\psi$—that skips verbal evaluations and directly maps inputs $(x,\mathcal{Y},f_{i},y)$ to ratings $r=\psi(x,\mathcal{Y},f_{i},y)$. Namely, any tuple $(x,\mathcal{Y},f_{i},e)$ can be converted into supervised training data $\big{(}(x,\mathcal{Y},f_{i},y),\phi(x,\mathcal{Y},f_{i},e,y)\big{)}_{y\in\mathcal{Y}}$ for $\psi$, containing simulated ratings $r=\phi(x,\mathcal{Y},f_{i},e,y)$. The hope is that the individual preference model $\psi$ would be able to simulate the rating of any evaluator (represented by their features $f_{i}$), as long as the evaluator, prompt, and response set come from the same distribution as the one $\psi$ was trained on. Similar to the preference models used in current RLHF, $\psi$ could finally be used to fine-tune the actual AI system or steer its behavior in real time. In fact, if the evaluators’ features $f_{i}$ are omitted in the training process sketched above, $\psi$ reduces to the standard preference model as is already used in RLHF. This is vulnerable to, for example, evaluators that strategically misreport (Siththaranjan et al., 2023), or to issues that arise from a disproportionate representation in the set of evaluators. We add evaluators’ features $f_{i}$ to $\psi$ so that preferences can later be aggregated in a transparent and deliberate manner via an additional social choice step, as we discuss next.
6 How Do We Incorporate Diverse Individual Feedback?
@@
Here we sketch several variants of two approaches for including diverse input or feedback into AI systems in a consistent way using methods from social choice theory. The first proposes adding an additional preference aggregation step somewhere during training, thereby turning RLHF into RLCHF: Reinforcement Learning from Collective Human Feedback. The second approach instead proposes adding an additional simulated collective decision step somewhere in the training or the system’s real-time decision procedure, similar to Bakker et al. (2022) and Jarrett et al. (2023).
@@ 6.1 Proposal: Reinforcement Learning from Collective Human Feedback (RLCHF)
Preference aggregation could be incorporated into RLHF in several ways, early on or rather late. For simplicity, assume a base version of rankings-based RLHF that (1) takes a database of prompts $x$ together with corresponding sets of possible responses $\mathcal{Y}$, (2) asks one associated evaluator $i(x,\mathcal{Y})$ to provide a ranking $R(x,\mathcal{Y})$ of the elements of $\mathcal{Y}$, (3) turns this ranking into $|\mathcal{Y}|$ many data points for training a common preference model $\varrho$ that produces numerical ratings $r=\varrho(x,y)$, and (4) uses these ratings as rewards in fine-tuning the actual LLM via reinforcement learning.
The earliest point one may introduce preference aggregation is between steps (2) and (3). Instead of a single evaluator, we may ask the members of a jury $J(x,\mathcal{Y})$ to provide individual rankings $R_{j}$. Using some ordinal social welfare function $F$, those rankings can then be aggregated into a collective ranking $R=F((R_{j})_{j\in J})$ to be used in step (3). We call this “RLCHF using aggregated rankings” (Fig. 2).
@@
Alternatively, one could use cardinal preference aggregation between steps (3) and (4). For this, change step (3) so that a model of individual preferences is trained, mapping pair $(x,\mathcal{Y})$ and evaluator $i$ with features $f_{i}$ to predicted ratings $r_{i}=\psi(x,f_{i},y)$. Also, generate a large sample of feature vectors $f_{1},\dots,f_{N}$ that is representative of the stakeholder population. Then a cardinal social welfare function $W$ can be used to aggregate simulated individual ratings into a social rating $\varrho(x,y)=W(\psi(x,f_{1},y),\dots,\psi(x,f_{N},y))$ which can be used in step (4). We call this “RLCHF using evaluator features and aggregated ratings” (Fig. 3).
@@ 6.2 Proposal: Simulated Collective Decisions
RLCHF, as described above, keeps the reinforcement learning step that requires numerical rewards, and it uses ordinal or cardinal preference aggregation to produce those rewards for all possible responses $y\in\mathcal{Y}$. A different approach would replace reinforcement learning by something else and introduce social choice methods in the form of simulated collective decisions rather than preference aggregation.
For one thing, one could modify “RLCHF using evaluator features and aggregated ratings” into “Supervised Learning from Simulated Collective Decisions”, as shown in Fig. 4. For this, in step (3) from above, use the individual preference model $r_{i}=\psi(x,f_{i},y)$ and feature vectors $f_{1},\dots,f_{N}$ not to produce an aggregated rating but to simulate a collective choice that picks a single winning response $y^{*}=C\big{(}(\psi(x,f_{j},y))_{y\in\mathcal{Y},j=1,\dots,N}\big{)}$. Here, $C$ is now a single-winner social choice function. Then in step (4), use data point $(x,y^{*})$ to train the actual AI system via supervised (rather than reinforcement) learning. Instead of picking a single winner $y^{*}$, we could also use a multi-winner social choice function $C$ that outputs, say, a set of three responses $(y^{\prime},y^{\prime\prime},y^{\prime\prime\prime})$. These can then be (creatively) combined into a single response, e.g., by merging them into a bullet-point list and adding “The following are (three) typical answers to your question: …” at the beginning.
A more radical modification would drop the fine-tuning-via-learning step altogether (leaving the LLM only pretrained) and instead simulate the collective choice at inference time. Whenever the live system is prompted with some $x$, generate $k\gg 1$ many candidate responses $y_{i}$ and $N\gg 1$ many evaluator feature vectors $f_{j}$ representative of the stakeholder population for the problem $(x,\mathcal{Y})$, and then directly return the winner $y^{*}=C\big{(}(\psi(x,f_{j},y_{i}))_{j,i=1}^{N,k}\big{)}$ of the simulated collective choice. Here, too, $C$ could be a multi-winner or probabilistic social choice rule.